# From Semantic To Instance: A Semi-Self-Supervised Learning Approach

### v1

@@ Abstract
Keywords: Instance Segmentation, Semi-Supervised Learning, Self-Supervised Learning, Wheat Head Instance Segmentation, Precision Agriculture, Data Synthesis
@@ 1 Introduction
Figure 1: Overview of the proposed approach. A semantic segmentation model is used to generate the semantic mask of the image, along with the grayscale and L channel of LAB color maps, which are concatenated to generate the new image-mask representation to feed to the instance segmentation model.
@@ 2 Related Works
In this context, numerous studies have explored the utility of wheat head semantic segmentation in top-view images of real fields. Rawat et al. [30] experimented with the wheat head semantic segmentation task with a pool-based active learning approach, obtaining an Intersection over Union (IoU) of $0.70$ on only the 2020 version of the UTokyo subset of the Global Wheat Head Detection (GWHD) dataset [7]. Najafian et al. [25] proposed a semi-self-supervised approach for wheat head segmentation, generating synthetic data from limited annotations and employing domain adaptation, resulting in a Dice score of $0.91$. Myers et al. [24] enhanced synthetic data realism by training a CycleGAN to transform unrealistic synthetic data into more authentic images, achieving a Dice score of $0.796$ on the GWHD dataset [7] after training the model on the GAN-generated data and final applying pseudo-learning phase. In addition, Ghanbari et al. [12] studied the need for domain adaptation and data diversity creating large-scale simulated data in training deep models, evaluated specifically on a wheat head segmentation task. Ennadifi et al. [9] developed a DeepMAC [9] and YOLOv5 [38] model for single-head segmentation of wheat heads, individually extracted using wheat head bounding boxes from $701$ manually annotated images or YOLOv5 model-generated bounding boxes, achieved an average F1 score of $86\%$ with their best-performing model.
@@
Li et al. [17] utilized transfer learning to fine-tune SAM model [16] for crop disease and pest segmentation in agricultural domains. They employed three datasets: two for coffee-leaf diseases, incorporating $1,100$ images, and one for pest image segmentation, containing $5,464$ images of $10$ common agricultural pests. The coffee-leaf disease datasets were split into training and testing subsets with an $80/20$ ratio. The pest dataset was divided into two groups, each with five pest categories. Their final model achieved an IoU of $59.58\%$ on the coffee-leaf disease datasets and $89.56\%$ on the pest datasets.
@@
With regard to the advancements in state-of-the-art (SOTA) instance segmentation models, such as the Segment Anything Model (SAM) series [16, 29], we opted to utilize YOLOv9 [39]—recognized as the SOTA in object detection and localization, known for its accuracy and efficiency in instance segmentation—in this work. SAM, which has outperformed models including FastSAM [43], was extensively trained on over $1.1$ billion segmentation masks across 11 million licensed images. FastSAM, on the other hand, leverages YOLOv8 [31] and a CLIP [29] prompting head, achieving 50×50$\times$50 × better speed but with reduced performance compared to SAM. SAM2 [29] introduced memory mechanisms for real-time image and video segmentation. However, YOLOv9 stands out by introducing the integration of Programmable Gradient Information (PGI) to address the information bottleneck problem and proposing the Generalized Efficient Layer Aggregation Network (GELAN) for optimal parameter utilization and computational efficiency. This architecture’s capacity to flexibly integrate various computational blocks while maintaining high performance makes YOLOv9 particularly adaptable to a wide range of applications, including instance segmentation, aligning better with our work’s objectives.
@@ 3.1 Data
For model evaluation, we manually annotated a set of $365$ images, hereafter referred to as $\textit{GHD}_{te}$. The average number of wheat head instances for these images was $52.3$, ranging from 00 to $102$. Also, a set of 100 images of size $1024\times 1024$, referred to as $\textit{LateStage}_{te}$, were used for further model evaluation. Each frame in $\textit{LateStage}_{te}$ has an average of $58.12$ instances, ranging from $5$ to $146$. Figure 3 presents a few manually annotated examples from our evaluation sets.
@@ 3.2 Data Synthesis
We adopted the cut-and-paste methodology originally proposed by [26] for object detection and applied it to the instance segmentation data synthesis with some modifications. As the primary adjustment, we increased the number of manually annotated frames—containing larger wheat head objects—to $10$, categorized as $\textit{W}_{tr}$ and $\textit{W}_{va}$. Additionally, we incorporated a wider range of video frames from various background scenes, sourced from $\textit{B}_{tr}$ and $\textit{B}_{va}$, to enhance the final synthesized data variability.
[…]
Figure 4: Data synthesis pipeline for instance segmentation, adapted from [25], extracts fake and real wheat heads from manually-annotated images (detailed in [25]), categorized into Harvest-ready ($\textit{W}_{tr}$) and Heading-complete ($\textit{W}_{va}$) stages subsets. Background frames were similarly extracted, diversified, and categorized into $\textit{B}_{tr}$ and $\textit{B}_{va}$. Fake ($\textit{F}_{tr}$ and $\textit{F}_{va}$) and real ($\textit{R}_{tr}$ and $\textit{R}_{va}$) wheat heads were overlaid on the backgrounds, with smaller heads used when overlaying more than $50$ objects and larger heads used for fewer objects, synthesizing field images captured at various altitudes with the corresponding instance segmentation maps.
@@
To synthesize each sample, we randomly selected $10$ to $100$ fake and then real wheat heads to overlay on the backgrounds. However, dissimilar to [25], we divided the real wheat heads into two subsets based on the object’s larger dimensions, with smaller and larger wheat heads in a $50\%-50\%$ ratio. Overlaying $10$ to $50$ objects, we selected from the subset containing larger heads; for more than $50$ overlying objects, we chose from the subset with smaller wheat heads. This approach aimed to enhance the realism of the synthesized samples by simulating wheat field images captured from varying altitudes. Additionally, we designed our pipeline to computationally generate the instance segmentation maps and contour labels, required for YOLO models, in parallel with synthesizing each image, for model development.
@@
For training with the YOLOv9 [39] model, we chose not to apply additional online augmentations, relying solely on the augmentations integrated into YOLOv9. Instead, we increased the variability of our dataset by applying strong augmentations to the synthesized samples, in an offline manner. Specifically, we applied Flip, Rotation, and Elastic transformations for fake and real wheat head objects before overlying them. Furthermore, we opted for a long list of pixel-level augmentations including Color Jitter, Channel Shuffle and Dropout, Solarize, Blur, and Noise for augmenting the synthesized images at the end of the process. We employed the Albumentations [2] Python package, version 1.4.14, in implementing all the augmentation transformations.
3.3 Rotation-Augmented Data
@@
For the second phase of the model training, we utilized real wheat field images to implement our domain adaptation method. As previously discussed, pixel-accurate instance annotation of many agricultural images is infeasible, and the $19$ samples in $\textit{Rotated}_{tr}$ were insufficient for fine-tuning a large model such as YOLOv9. Addressing these limitations, we extended the $\textit{Rotated}_{tr}$ dataset by applying image rotation. To enhance variability in our limited training dataset and better represent natural environments, we utilized rotation to simulate various angled top-down views of bending wheat stems, particularly effective for aerial imaging in windy conditions. Following the method used by [25], each image/mask pair was rotated with 00 to $259$ degrees, resulting in a large-scale dataset of $6840$ samples. To maintain the authenticity of samples for addressing domain shifts between synthetic and real data, we chose to exclude additional pixel-level or spatial-level transformations, apart from those applied by the YOLOv9 models in an online manner. We repeated the same process for the $\textit{Rotated}_{va}$ dataset, expanding it to match the sample size of $\textit{Rotated}_{tr}$.
@@
3.4 GLMask Generation
We utilized the Grayscale (G) and the perceptual lightness channel of LAB (L) color space versions of each RGB image, along with the corresponding semantic mask (M), to create a new image-mask representation of the RGB image, referred to as GLMask. We used pseudo-labels generated by the model developed for wheat head semantic segmentation by Najafian et al. [25]. By embedding the semantic mask, we aimed at efficient segmentation transfer from semantic to instance masks. Additionally, we designed GLMask to address a common issue in convolutional neural networks, where models often become overly reliant on color features. This reliance can hinder model generalization in the context of a diverse wheat dataset, as wheat color varies significantly due to growth stages and changing outdoor lighting conditions. By incorporating grayscale and the L channel of the LAB color spaces, GLMask intentionally reduces color information, encouraging the model to prioritize shape, texture, and pattern features.
3.5 Pseudo-Labeling
We explored pseudo-labeling as an alternative to rotation augmentation utilized for domain adaptation. This widely used self-supervised approach enhances model generalization by leveraging larger pseudo-labeled datasets, improving performance and reducing overfitting, while also lowering labeling costs, particularly in domains such as precision agriculture with packed objects. We utilized the best-performing model in the first phase, trained on the synthetic data, to label the unannotated $\textit{Pseudo}_{tr}$ and  $\textit{Pseudo}_{va}$ sets. Subsequently, we used these sets to fine-tune the model, initially trained on only synthetic data. We compare the performance of this model with the model developed using $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$ datasets to obtain our best-performing model as a better domain adaptation technique.
@@ 3.6 Experiments
As described in section 2, we selected the pretrained extended model YOLOv9-Seg architecture with $60,512,800$ parameters for all our experiments. We trained and evaluated the model using the data outlined in Table 1, with training conducted for $100$ epochs in each training phase on $1024\times 1024$ images. Using the AdamW optimizer [20], we set an initial learning rate of $31e-$ 3 to train the model on the $\textit{SYN}_{tr}$ and $\textit{SYN}_{va}$ datasets, resulting in SynModel, and a learning rate of $41e-$ 4 to fine-tune the $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$ datasets, leading to the development of RoAModel. We also applied a learning rate of $41e-$ 4 and used the $\textit{Pseudo}_{tr}$ and $\textit{Pseudo}_{va}$ datasets to conduct the pseudo-labeling experiment by fine-tuning model SynModel in comparison to model RoAModel, yielding PseModel. We further developed a baseline model, BaseModel, identical to SynModel but on RGB images. The batch size was set to $18$ and the IoU to $0.7$. The deterministic and seed parameters were set to ensure reproducibility, and the remaining parameters were kept at their default values. The default data augmentations applied included HSV color augmentation, horizontal flip, mosaic augmentation, erasing, translation, and scaling.
We subsequently trained two models on the Microsoft COCO 2017 dataset, using its training and validation sets, with the same exact parameters as those utilized in the wheat experiments with some exceptions. These models were trained on both the RGB and GLM versions of the dataset, with an adjusted batch size of $48$, enabled by a training image size of $640\times 640$ (up from the batch size of $18$ for $1024\times 1024$ images in wheat experiments). Given the dataset’s substantial size, both COCO models were trained for $50$ epochs. Additionally, we set IoU and conf parameters to be 0.6 and 0.25, respectively, during the evaluation. These two parameters enhanced both RGB and GLM model performances on the COCO validation set. We evaluate the models on the validation set of this dataset, as annotations for its test subset were unavailable. Note that we only require the binary (foreground vs. background) mask to create the GLM version of the data. This implies that, unlike the binary mask used as the semantic mask for the single-class wheat dataset, it serves as a binary mask instead of the semantic mask for the COCO dataset, which contains 80 distinct object classes.
All experiments were executed on three GPUs, specifically two A40 GPUs each equipped with $48$ GB GDDR6, and one Tesla V100S PCIe with $32$GB GPU. The environment was configured with Python 3.12, Ultralytics version 8.2.79, and PyTorch 2.4.0+cu124.

### v2

@@ Abstract
keywords:
Instance Segmentation , Semi-Supervised Learning , Self-Supervised Learning , Wheat Head Instance Segmentation , Precision Agriculture , Data Synthesis
[…]
@@ 1 Introduction
Figure 1: Schematic overview of the proposed Semi-Self-Supervised Learning Framework. The methodology proceeds in three logical stages: (Left) GLMask Representation: To reduce color dependency and enforce structural learning, the input RGB image is decomposed into Grayscale and L-channel (of CIELAB) components, then concatenated with a Semantic Segmentation Mask prior. (Right) Two-Stage Training Strategy: The framework proceeds with (1) Synthetic Pre-training, where a YOLOv9 model is trained on a large-scale synthetic dataset generated via a cut-and-paste pipeline (Sec 3.3); followed by (2) Domain Adaptation, where the model weights are transferred (horizontal dotted arrow) and fine-tuned on a rotation-augmented real dataset to bridge the domain gap (Sec 3.4).
@@ 2 Related Works
In this context, numerous studies have explored the utility of wheat head semantic segmentation in top-view images of real fields. Rawat et al. [30] experimented with the wheat head semantic segmentation task with a pool-based active learning approach, obtaining an Intersection over Union (IoU) of $0.70$ on only the 2020 version of the UTokyo subset of the Global Wheat Head Detection (GWHD) dataset [7]. Najafian et al. [25] proposed a semi-self-supervised approach for wheat head segmentation, generating synthetic data from limited annotations and employing domain adaptation, resulting in a Dice score of $0.91$. Myers et al. [24] enhanced synthetic data realism by training a CycleGAN to transform unrealistic synthetic data into more authentic images, achieving a Dice score of $0.796$ on the GWHD dataset [7] after training the model on the GAN-generated data and finally applying a pseudo-learning phase. In addition, Ghanbari et al. [12] studied the need for domain adaptation and data diversity, creating large-scale simulated data in training deep models, evaluated specifically on a wheat head segmentation task. Ennadifi et al. [9] developed a DeepMAC [9] and YOLOv5 [38] model for single-head segmentation of wheat heads, individually extracted using wheat head bounding boxes from $701$ manually annotated images or YOLOv5 model-generated bounding boxes, achieved an average F1 score of $86\%$ with their best-performing model.
@@
Li et al. [17] utilized transfer learning to fine-tune the SAM model [16] for crop disease and pest segmentation in agricultural domains. They employed three datasets: two for coffee-leaf diseases, incorporating $1,100$ images, and one for pest image segmentation, containing $5,464$ images of $10$ common agricultural pests. The coffee-leaf disease datasets were split into training and testing subsets with a $80/20$ ratio. The pest dataset was divided into two groups, each with five pest categories. Their final model achieved an IoU of $59.58\%$ on the coffee-leaf disease datasets and $89.56\%$ on the pest datasets.
@@
With regard to the advancements in state-of-the-art (SOTA) instance segmentation models, such as the Segment Anything Model (SAM) series [16, 29], we opted to utilize YOLOv9 [39]—recognized as the SOTA in object detection and localization, known for its accuracy and efficiency in instance segmentation—in this work. SAM, which has outperformed models including FastSAM [43], was extensively trained on over $1.1$ billion segmentation masks across 11 million licensed images. FastSAM, on the other hand, leverages YOLOv8 [31] and a CLIP [29] prompting head, achieving $50\times$ better speed but with reduced performance compared to SAM. SAM2 [29] introduced memory mechanisms for real-time image and video segmentation. However, YOLOv9 stands out by introducing the integration of Programmable Gradient Information (PGI) to address the information bottleneck problem and proposing the Generalized Efficient Layer Aggregation Network (GELAN) for optimal parameter utilization and computational efficiency. This architecture’s capacity to flexibly integrate various computational blocks while maintaining high performance makes YOLOv9 particularly adaptable to a wide range of applications, including instance segmentation, aligning better with our work’s objectives.
@@ 3 Material and Methods
We propose a Semi-Self-Supervised learning framework designed to segment densely packed objects with minimal manual annotation. As illustrated in Figure 1, our approach operates in three stages: (1) Representation Learning, where we introduce GLMask to shift the model’s focus from color to structure; (2) Synthetic Pre-training, utilizing a cut-and-paste synthesis pipeline to initialize weights; and (3) Domain Adaptation, employing a rotation-based consistency strategy to bridge the gap between synthetic and real-world data. This pipeline enables the transition from coarse semantic signals to precise instance masks. To operationalize this framework, specifically the data synthesis and generation phases, a curated set of minimal real-world samples is required. We therefore begin by detailing these fundamental data requirements.
@@ 3.1 Data
For model evaluation, we manually annotated a set of $365$ images, hereafter referred to as $\textit{GHD}_{te}$. The average number of wheat head instances for these images was $52.3$, ranging from 0 to $102$. Also, a set of 100 images of size $1024\times 1024$, referred to as $\textit{LateStage}_{te}$, were used for further model evaluation. Each frame in $\textit{LateStage}_{te}$ has an average of $58.12$ instances, ranging from $5$ to $146$. Figure 3 presents a few manually annotated examples from our evaluation sets.
@@
3.2 GLMask Representation
As the foundational component of our pipeline, we design an input format that explicitly leverages semantic priors to guide the transformation from semantic to instance-level segmentation. Specifically, we utilized the Grayscale (G) and the perceptual lightness channel of LAB (L) color space versions of each RGB image, along with the corresponding semantic mask (Mask), to create a new image-mask representation of the RGB image, referred to as GLMask.
For the semantic mask component, we used pseudo-masks generated by the model developed for wheat head semantic segmentation by Najafian et al. [25], which reportedly achieves a Dice score of 0.91 on wheat datasets. By embedding the semantic mask, we aimed at efficient segmentation transfer from semantic to instance masks. Additionally, we designed GLMask to address a common issue in convolutional neural networks, where models often become overly reliant on color features. This reliance can hinder model generalization in the context of a diverse wheat dataset, as wheat color varies significantly due to growth stages and changing outdoor lighting conditions. By incorporating grayscale and the L channel of the LAB color spaces, GLMask intentionally reduces color information, encouraging the model to prioritize shape, texture, and pattern features.
The GLMask representation was specifically designed as a three-channel input (G, L, M) to address the input requirements of the chosen pre-trained YOLOv9e-Seg architecture while ensuring the model learns features based on shape and density rather than fluctuating spectral signatures. This configuration prioritizes domain-invariant structural information over raw color, which often varies across different growth stages and environments. Consequently, this design choice targets the most informative, color-free components as a principled design constraint, focusing the model’s capacity on features that persist across the diverse wheat domains.
@@ 3.3 Data Synthesis
To effectively train our model using the proposed GLMask representation without relying on extensive manual annotation, we leveraged a synthetic data generation strategy. We adopted the cut-and-paste methodology originally proposed by [26] for object detection and applied it to the instance segmentation data synthesis with some modifications. As the primary adjustment, we increased the number of manually annotated frames—containing larger wheat head objects—to $10$, categorized as $\textit{W}_{tr}$ and $\textit{W}_{va}$. Additionally, we incorporated a wider range of video frames from various background scenes, sourced from $\textit{B}_{tr}$ and $\textit{B}_{va}$, to enhance the final synthesized data variability.
In line with the methodology described by [26, 25], we extracted two sets of fake ($\textit{F}_{tr}$ and $\textit{F}_{va}$) and real wheat head objects ($\textit{R}_{tr}$ and $\textit{R}_{va}$), which were randomly overlaid on the background images in order. We used $\textit{B}_{tr}$$\textit{F}_{tr}$ and $\textit{R}_{tr}$ for large-scale training data synthesis of $20,000$ sample size, hereby denoted as $\textit{SYN}_{tr}$. Additionally, we employed $\textit{B}_{va}$$\textit{F}_{va}$ and $\textit{R}_{va}$ in synthesizing a validation set of size $10,000$, entitled $\textit{SYN}_{va}$, with a different distribution from the training set. The data synthesis pipeline is visually demonstrated in Figure 4.
Figure 4: Detailed workflow of the data synthesis pipeline for instance segmentation, adapted from [25]. The process is strictly partitioned into Train Split (left, yellow) and Valid Split (right, green) streams to ensure distinct data distributions, utilizing wheat heads categorized into Harvest-ready ($\textit{W}_{tr}$) and Heading-complete ($\textit{W}_{va}$) stages along with their corresponding diversified background frames ($\textit{B}_{tr}$ and $\textit{B}_{va}$). The pipeline integrates three source components: background frames ($B$), synthetic fake foregrounds ($F$), and extracted real foregrounds ($R$). The synthesis proceeds sequentially: (1) Fake Object Overlay: Backgrounds are first populated with synthetic wheat templates to create a dense underlying texture; (2) Real Object Overlay & Mask Generation: Real wheat instances are superimposed using density-dependent selection logic (using smaller heads in blue dashed boxes for $\geq 50$ objects and larger heads in red dashed boxes for fewer objects) to simulate varying altitudes, while simultaneously generating pixel-perfect instance masks; and (3) Data Augmentation: The composited images undergo photometric and geometric transformations to produce the final synthetic training ($\textit{SYN}_{tr}$) and validation ($\textit{SYN}_{va}$) datasets.
@@
To synthesize each sample, we randomly selected $10$ to $100$ fake and then real wheat heads to overlay on the backgrounds. However, dissimilar to [25], we divided the real wheat heads into two subsets based on the object’s larger dimensions, with smaller and larger wheat heads in a $50\%-50\%$ ratio. Overlaying $10$ to $50$ objects, we selected from the subset containing larger heads; for more than $50$ overlying objects, we chose from the subset with smaller wheat heads. This approach aimed to enhance the realism of the synthesized samples by simulating wheat field images captured from varying altitudes. Additionally, we designed our pipeline to computationally generate the instance segmentation maps and contour labels required for YOLO models, in parallel with synthesizing each image, for model development.
@@
For training with the YOLOv9 [39] model, we chose not to apply additional online augmentations, relying solely on the augmentations integrated into YOLOv9. Instead, we increased the variability of our dataset by applying strong augmentations to the synthesized samples, in an offline manner. Specifically, we applied Flip, Rotation, and Elastic transformations for fake and real wheat head objects before overlaying them. Furthermore, we opted for a long list of pixel-level augmentations, including Color Jitter, Channel Shuffle and Dropout, Solarize, Blur, and Noise for augmenting the synthesized images at the end of the process. We employed the Albumentations [2] Python package, version 1.4.14, in implementing all the augmentation transformations.
3.4 Domain Adaptation via Rotation-Augmented Data
@@
To bridge the inevitable domain shift between the synthetic data generated in the previous step and real-world agricultural environments, we prepared a real-world dataset for domain adaptation. For this second phase of the model training, we utilized real wheat field images to implement our domain adaptation method. As previously discussed, pixel-accurate instance annotation of many agricultural images is infeasible, and the $19$ samples in $\textit{Rotated}_{tr}$ were insufficient for fine-tuning a large model such as YOLOv9. Addressing these limitations, we extended the $\textit{Rotated}_{tr}$ dataset by applying image rotation. To enhance variability in our limited training dataset and better represent natural environments, we utilized rotation to simulate various angled top-down views of bending wheat stems, particularly effective for aerial imaging in windy conditions. Following the method used by [25], each image/mask pair was rotated with 0 to $259$ degrees, resulting in a large-scale dataset of $6840$ samples. To maintain the authenticity of samples for addressing domain shifts between synthetic and real data, we chose to exclude additional pixel-level or spatial-level transformations, apart from those applied by the YOLOv9 models in an online manner. We repeated the same process for the $\textit{Rotated}_{va}$ dataset, expanding it to match the sample size of $\textit{Rotated}_{tr}$.
@@
3.5 Domain Adaptation via Pseudo-Labeling
To evaluate the efficacy of our primary rotation-based adaptation strategy, we investigated Pseudo-Labeling as a comparative self-supervised baseline. We explored this widely used approach as an alternative for domain adaptation, as it is known to enhance model generalization by leveraging larger pseudo-labeled datasets, improving performance and reducing overfitting while lowering labeling costs, particularly in domains such as precision agriculture with packed objects. We utilized the best-performing model in the first phase, trained on the synthetic data, to label the unannotated $\textit{Pseudo}_{tr}$ and  $\textit{Pseudo}_{va}$ sets. Subsequently, we used these sets to fine-tune the model, initially trained on only synthetic data. We compare the performance of this model with the model developed using $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$ datasets to obtain our best-performing model as a better domain adaptation technique.
@@ 3.6 Experiments
As described in section 2, we selected the pretrained extended model YOLOv9e-Seg architecture with $60,512,800$ parameters for all our experiments. We trained and evaluated the model using the data outlined in Table 1, with training conducted for $100$ epochs in each training phase on $1024\times 1024$ images. Using the AdamW optimizer [20], we set an initial learning rate of 1e31e-3 to train the model on the $\textit{SYN}_{tr}$ and $\textit{SYN}_{va}$ datasets, resulting in SynModel, and a learning rate of 1e41e-4 to fine-tune the $\textit{Rotated}_{tr}$ and $\textit{Rotated}_{va}$ datasets, leading to the development of RoAModel. We also applied a learning rate of 1e41e-4 and used the $\textit{Pseudo}_{tr}$ and $\textit{Pseudo}_{va}$ datasets to conduct the pseudo-labeling experiment by fine-tuning model SynModel in comparison to model RoAModel, yielding PseModel. We further developed a baseline model, BaseModel, identical to SynModel but on RGB images. The batch size was set to $18$ and the IoU to $0.7$. The deterministic and seed parameters were set to ensure reproducibility, and the remaining parameters were kept at their default values. The default data augmentations applied included HSV color augmentation, horizontal flip, mosaic augmentation, erasing, translation, and scaling.
We subsequently trained two models on the training split of the Microsoft COCO 2017 dataset, with the same exact parameters as those utilized in the wheat experiments with some exceptions. We utilized the validation split (5k images) exclusively for evaluation to ensure zero data leakage during training. These models were trained on both the RGB and GLMask versions of the dataset, with an adjusted batch size of $48$, enabled by a training image size of $640\times 640$ (up from the batch size of $18$ for $1024\times 1024$ images in wheat experiments). Given the dataset’s substantial size, both COCO models were trained for $50$ epochs. Additionally, we set IoU and conf parameters to be 0.6 and 0.25, respectively, during the evaluation. These two parameters enhanced both RGB and GLMask model performances on the COCO validation set. We evaluate the models on the validation set of this dataset, as annotations for its test subset were unavailable.
Note that we only require the binary (foreground vs. background) mask to create the GLMask version of the data. These masks were generated by aggregating all ground-truth instance annotations into a single binary foreground channel. This implies that, unlike the binary mask used as the semantic mask for the single-class wheat dataset, it serves as a binary mask instead of the semantic mask for the COCO dataset, which contains 80 distinct object classes.
All experiments were executed on three GPUs, specifically two A40 GPUs each equipped with $48$ GB GDDR6, and one Tesla V100S PCIe with 3232GB GPU. The environment was configured with Python 3.12, Ultralytics version 8.2.79, and PyTorch 2.4.0+cu124.